A man is 30 years old and his son is 13 years old. In how many years will the son be 3/4s as old as his father?: A man is 30 years old and his son is 13 years old. In how many years will the son be 3/4s as old as his father?
Answer by mangopeeler07(429)

man=30
son=13
x=amount of years

13+x=(3/4)(30+x)

Distribute
13+x=22.5+3/4x

Subtract 13 from both sides
x=9.5+3/4x

Subtract 3/4x
1/4x=9.5

Multiply both sides by 4
x=38

38 years